Cos-1(12/13)+2cos-1(√64/√65)+cos-1(√49/√50)=cos-1(1/√2)

Grade:12

1 Answers

Vikas TU
4 years ago
Dear stuedent 
Cos-1(12/13) = tan-1 (5/12)
cos-1(√64/√65) = tan-1 (1/8) 
cos-1(√49/√50) = tan-1 (1/7)
Now use tan-1 (a) + tan-1(b) formula and find a tan-1(c) 
Now again find tan-1(c) + tan-1(d) and equate 
LHS = RHS 
Hope this helps 
Good luck